OpenStudy (anonymous):

What are the values of x and y in this linear system? 2x + 3y = 8 x + 2y = 12

OpenStudy (anonymous):

A. x = 24, y = 8 B. x = 16, y = -20 C. x = -8, y = -16 D. x = -20, y = 16

OpenStudy (anonymous):

@gorv @g17

OpenStudy (anonymous):

Well what method have you been doing lately to solve this type of problem?

OpenStudy (anonymous):

Solving Linear Systems Using Substitution

OpenStudy (anonymous):

x+2y=12 x=12-2y put this value of x in other equation and solve

OpenStudy (anonymous):

Well ok the best way I see is if you solve for x on both equation since the bottom equation already has x by itself with no coefficient.

OpenStudy (anonymous):

Or you don't have to solve for x on the top equation sorry....thats just the way I like to do it just follow what g17 does it seems to be easier to follow.

OpenStudy (anonymous):

2(12-2y)+3y=8?

OpenStudy (anonymous):

Yes now distribute and solve for y

OpenStudy (anonymous):

24-4y+3y=8 24 -1y = 8 -24 -24 ------------- -1y=-16 y= 16?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yeah now all you need is the x.....just play y=16 to either equation and solve for x again. and you are done. good job